OpenStudy (anonymous):

find the value for y for this set 12x-5y=30 y=2x-6

OpenStudy (saifoo.khan):

x = 0 y = -6

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

dang you are quick

OpenStudy (katrinakaif):

Plug it into the orginal equation 12 x - 5(2x - 6) = 30 Distribute 12x - 10x +30 = 30 Collect like terms 2x = 0 x = 0 Plug itno the y equation: y = 2(0) - 6 y = 0 - 6 y = -6

OpenStudy (saifoo.khan):

:D